OpenStudy (carolina4567):

In △PQR, find the measure of ∡P. Triangle PQR where angle Q is a right angle. QR measures 33 point 8. PR measures 57 point 6. Measure of angle P is unknown. 30.4° 35.9° 59.6° 54.1°

OpenStudy (carolina4567):

I need help with this question asap. Please help

OpenStudy (anonymous):

the best thing to do would be to draw a picture first

OpenStudy (carolina4567):

I have the picture one sec let me see how i can post it

OpenStudy (carolina4567):

Its more like this

OpenStudy (anonymous):

oh PR is the hypontenuse use \[\theta=\sin ^{-1}\left( \frac{ 33.8 }{ 57.6 } \right)\]

OpenStudy (anonymous):

sin = opposite/hypontenuse

OpenStudy (carolina4567):

Ok so i need to solve

OpenStudy (anonymous):

and inverse sin will give you the angle

OpenStudy (anonymous):

yes

OpenStudy (carolina4567):

Ok thank you i get it now thank you

OpenStudy (anonymous):

you're welcome